OpenStudy (anonymous):

Jeff works as a pharmacuetical representative. He receives a 6% monthly commission on sales up to $60,000. He receives an 8.5 commission on all sales above $60,000. If he sold $86,000 worth of his product one month, how much did he receive in commission? Please help.

OpenStudy (anonymous):

60,000*6% + 26,000* 8.5% = 3600+ 2210 = 5810